The velocity of particle moving along x axis is given T>0 by Vx=(32.0t-2.00t3)m/s. Where t is in s. What is the acceleration of the particle when (after t=0) it achieves its maximum displacement in the positive x direction?


1
Expert's answer
2022-01-26T07:40:20-0500

To find the time at which the displacement is the maximum, equate velocity to zero. Then find the derivative of velocity t get the acceleration.


"v_x=30t-2t^3=0,\\
t=3.87\text{ s},\\
a_x(t)=30-6t^2→a(3.87)=15\text{ m/s}^2."